OpenStudy (anonymous):

how do i solve x^2 + 5x + 6 = 0

OpenStudy (anonymous):

Is this how the question is phrased ?

OpenStudy (unklerhaukus):

\[x^2 + 5x + 6 = 0\] \[x^2 + 2x+3x + 6 = 0\] \[(x+3)(x+2)=0\]

OpenStudy (anonymous):

\[x^{2} + 5x + 6 = 0\]

OpenStudy (anonymous):

exactly what @UnkleRhaukus is doing. you factor it and then put both of them equal to 0

OpenStudy (anonymous):

so (x+3)=0 and (x+2)=0 and solve for x and you get....??

OpenStudy (anonymous):

i figure it out thank you

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

x2+5x+6=0 you're going to factor. what you do is try to figure out what 2 numbers multiply to give you 6 but add to give you 5. In this case you would have 3 & 2 x2+2x+3x+6=0 Group them together (x2+2x)+(3x+6)=0 Factor out what ever is possible here its the x in the first group & 3 in the second x(x+2)+3(x+2)=0 from here you set both equations equal to 0 an solve for x (x+2)=0 x=-2 (x+3)=0 x=-3 and that is your answer x=-2 & x=-3